OpenStudy (anonymous):

3x-5y=9 when x= -1,0,1

OpenStudy (anonymous):

Plug in X 3 x -1 - 5y = 9 -3 - 5y = 9 Add -3 to both sides: 5y = 6 Divide: y = 1.2 3 x 0 - 5y = 9 0 - 5y = 9 Add 0 to both sides: 5y = 9 Divide: y = 1.8 3 x 1 - 5y = 9 3 - 5y = 9 Add 3 to both sides: 5y = 12 Divide: y = 2.4